OpenStudy (larseighner):

(\ f(x) = \sqrt{x+1} \)

OpenStudy (larseighner):

(/ f(x) = \sqrt{x+1} /)

OpenStudy (larseighner):

/( f(x) = \sqrt{x+1} /)

OpenStudy (larseighner):

\( f(x) = \sqrt{x+1} \)

OpenStudy (abhisar):

``` \(\huge f(x) = \sqrt{x+1}\) ``` \(\huge f(x) = \sqrt{x+1}\)

OpenStudy (abhisar):

Hello @LarsEighner ! U need help ?

OpenStudy (larseighner):

\( {\displaystyle \color{red}{\lim_{h \rightarrow 0} { {f(x+h) - f(x)} \over h} }} \)

OpenStudy (abhisar):

\({\huge\displaystyle \color{green}{\lim_{h \rightarrow 0} { {f(x+h) - f(x)} \over h} }}\)

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (larseighner):

@abhisar Thanks I think I'm getting it. I am a mathjax wizard but I just need to find the knobs hers.

OpenStudy (abhisar):

oh i thought u need help

OpenStudy (abhisar):

\(\huge\color{green}{ت}\)

OpenStudy (abhisar):

"I just need to find the knobs hers"......sorry i didn't got this !

OpenStudy (larseighner):

\( \begin{align} f^\prime (3) &= \lim_{ h \rightarrow0} {{f(3 + h) -f(3)} \over { h}} \cr &= \lim_{ h \rightarrow0} {{(4+h)^{1\over 2} -2} \over { h}} \cr &= \lim_{ h \rightarrow0} {{[(4+h)^{1\over 2} -2][(4+h)^{1\over 2} +2]} \over { h[(4+h)^{1\over 2} +2]}} \cr &= \lim_{ h \rightarrow0} {{[(4+h)^{1\over 2}]^2 -4} \over { h[(4+h)^{1\over 2} +2]}} \cr &= \lim_{ h \rightarrow0} {{(4+h) -4} \over { h[(4+h)^{1\over 2} +2]}} \cr &= \lim_{ h \rightarrow0} {h \over { h[(4+h)^{1\over 2} +2]}} \cr &= \lim_{ h \rightarrow0} {1 \over { (4+h)^{1\over 2} +2}} \cr &= {1 \over { (4+0)^{1\over 2} +2}} \cr f^\prime (3) &= {1 \over 4} \end{align} \)

OpenStudy (larseighner):

hers = here

OpenStudy (abhisar):

oh !